ROTATIONAL AND FINE STRUCTURE ANALYSES OF THE HIGH-RESOLUTION LIF SPECTRA OF THE DEUTERATED ISOTOPOMERS OF THE METHOXY RADICAL

\underline{JINJUN LIU}, JOHN T. YI, VADIM STAKHURSKY and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210; DMITRY MELNIK and ROBERT F. CURL, Department of Chemistry and Rice Quantum Institute, Rice University, Houston, TX 77005.

The methoxy radical, CH_{3}O, is one of the most interesting and most widely-studied free radicals. It provides a benchmark case of the Jahn-Teller effect in a molecule with spin-orbit interaction. The partial deuteration of the methoxy radical breaks the molecular symmetry, thereby reducing the electronic orbital angular momentum in the X̃^{2}E ground state by introducing new terms in the rovibronic Hamiltonian of CH_{3}O. The resolved rotational and fine structure in the LIF spectra of CHD_{2}O and CH_{2}DO has been assigned and analyzed by simulating and fitting the 3_{0}^{2} and (6′)_{0}^{1} bands of the Ã^{2}A_{1}–X̃^{2}E_{3/2} transition for each isotopomer, providing molecular constants and energy levels for both states of each isotopomer. The ground state results were used to predict the previously observed X̃ state microwave spectrum.^{a} This prediction is consistent within the experimental accuracy of the LIF experiment. As an additional control, the 3_{0}^{2} and 6_{0}^{1} bands of the Ã^{2}A_{1}–X̃^{2}E_{3/2} transition for CH_{3}O and CD_{3}O were analyzed using the same method.^{b} Efforts are underway to fit simultaneously the LIF, microwave, and possibly SEP data of the ground state of these isotopomers to a rotational and fine structure Hamiltonian.

aD. Melnik, V. Stakhursky, V. A. Lozovsky, T. A. Miller, C. B. Moore and F. C. De Lucia, WJ09, 59^{th} International Symposium on Molecular Spectroscopy, 2004.

bY. Endo, S. Saito and E. Hirota, J. Chem. Phys. \underline{81}, 122 (1984).